2/3(7/8+4x)-5/8=3/8
Answer by nerdybill(7384)

2/3(7/8+4x)-5/8=3/8
.
Start by multiplying both sides by 24:
16(7/8+4x)-15=9
Distribute the 16 to terms inside the parenthesis on the left:
(14+64x)-15=9
14+64x-15=9
64x-1=9
64x=10
x = 10/64
x = 5/32